Reports 1-1 of 1 Clear search Modify search
AdV-FLT (Design/Construction)
bertolini, guo, vacatello - 13:41 Wednesday 05 November 2025 (68099) Print this report
Measuring vertical-to-horizontal coupling in the FCEM payload

Yesterday during the maintainance, we had another attempts to meausre the V-H coupling with the filter cavity end mirrors. 

Since during our last measurement, we noticed the coupling between the longitudinal and TX, which might be caused by the unbalanced actuators on the mirror, we first checked if this coupling could be reduced by changing the actuation matrix. Currently the Z direction of the mirror use only the top and bottom actuators, so we tried to adjust coeffcients between them. The following three combinations were tested for the [Top, bottom], [+1.078,-0.922] (starting point), [+1,-1], [+0.9,-1.1], with a line injection at 7Hz to the logitudinal loop to check the coupling. In Fig 1, the spectrum of the AA TX (indication of the coupling) zoomed around 7Hz is shown for three matices as blue, purple and orange. It seems with the driving of  [+1,-1], we reached a state with less coupling. Further test could be done to find the optimal actuation to further minimize the coupling. 

Then we repeated the line injection in low frequency to the suspension Y, but after each injection, we realigned cavity with the AA loops. The time for each injections are listed below, the amplitude of the lines are 30. 

4Hz  10:27:55 -- 10:31:00 UTC

5Hz   10:32:25 -- 10:35:20 UTC

7Hz   10:37:30 -- 10:39:50 UTC

Then we did a wide band noise injection between 7Hz and 50Hz, 

Filter cavity longitudinal lock open loop transfer function, noise amplitude of 3e-3: 10:46:25 -- 10:50:20 UTC

Noise injectio to FCEM suspension Y, noise amplitude 20: 11:05:15 -- 11:12:50 UTC

The analysis of the data will follow.  

Images attached to this report
Comments to this report:
bertolini, guo, vacatello - 8:02 Thursday 27 November 2025 (68260) Print this report

Data analysis from broadband noise injection

First we have measured the FC length control loop gain by injecting 7-50Hz white noise in the LFC control; the 3e-2 amplitude noise injection was performed from 2025-11-04-10h46m22-UTC to 2025-11-04-10h50m23-UTC. The measured G(f) open-loop TF is shown in the first plot:

We then proceeded to measure the V-H coupling by injecting noise into the vertical actuator of the FCEM GAS filter. A summary of the 7-50Hz noise injection is shown in the following plot which includes the spectrum of the cavity length error signal, the FCEM vertical actuation signal, which is DC calibrated, and the coherence. The peak at 26.55Hz in the cavity length error signal corresponds to the vertical bounce mode of the mirror suspension.

By combining the actuation signal with the SUMCON model of the payload mechanics we have estimated the actual vertical displacement of the mirror; the conversion also included the compensation of the 35-Hz pole in the coil driver response. The open-loop cavity length change dL can be estimated as dL = 5e-13*LFC_GR_PD_RF_5MHZ_I_CAL * (1+G(f)), where the scaling 1Hz = 0.5 pm is applied. From the comparison between the two reconstructed signals, see next plot, we can extract a V-H coupling of 2e-4 which appears to be frequency independent in the explored band. This measurement was taken with the FCEM mirror Z-actuation not optimized in terms of balancing; for this reason we are planning to repeat it under better balancing conditions.

 

 

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×